OpenStudy (anonymous):

Find the solution to the equation 64(3 – x) = 42x 64^(3 – x) = 4^(2x) 64 = 4^3 4^(3(3 – x)) = 4^2x 3(3 – x) = 2x 9 - 3x = 2x 9 = 5x x =9/5 I know how to solve it but what mehod am I using? and how would I explain it?

OpenStudy (compassionate):

Can you write the equation using the equation editor at the bottom left to make it easier to read? Thanks.

OpenStudy (anonymous):

Hold on I'll rewrite it so you can check that it's what was intended:

OpenStudy (anonymous):

thanks neil :)

OpenStudy (anonymous):

From here? \[64(3 – x) = 42x\] Or this? \[64^{3 – x} = 4^{2x}\]

OpenStudy (anonymous):

the latter.

OpenStudy (compassionate):

So top or bottom?

OpenStudy (anonymous):

i just answered the bottom one.

OpenStudy (anonymous):

9/5 is the real root for that problem, correct.

OpenStudy (anonymous):

That's the answer Neil, I need to know the name of the method I used to get it.

OpenStudy (compassionate):

Divide by whatever is multiplied in the squared terms.

OpenStudy (anonymous):

I would do the following first, both sides of the equations in terms of base 2

OpenStudy (anonymous):

Why base 2? because then you can take log\(_2\) ( ) to both sides ;-)

OpenStudy (anonymous):

This question is asking you to remember rules of exponents

OpenStudy (anonymous):

2^6 = 64 2^4 = 4

OpenStudy (anonymous):

so... \[\large 64^{3 – x} = 4^{2x}\] \[\large 2^{6(3 – x)} = 2^{2(2x)}\] \[\large \log_2 ( 2^{6(3 – x)} ) = \log_2 ( 2^{2(2x)})\] \[\large 6(3 – x) = 2(2x)\] \[\large 18 – 6x = 4x\] \[\large 18 = 10x\] 18/10 = 9/5

OpenStudy (anonymous):

Make sense @strcaitlyn13 ? :-D

OpenStudy (anonymous):

Yes i knew how to get the answer, just not what the method was.

OpenStudy (anonymous):

I like that you recognized that 64 and 4 were related by even factors. Alternately you could put both of them in terms of base 4 and take log\(_4\) to both sides instead. Works the same.

OpenStudy (anonymous):

But you need to use a log here I think or you're making an assumption that's false for some numbers, because the domain of logarithms are restricted from 0<x<\(\infty\)

OpenStudy (anonymous):

Minor detail, but some mathematicians and software engineers would make a big deal about it.

OpenStudy (anonymous):

i got it now thank you :)

OpenStudy (anonymous):

^_^